OpenStudy (loser66):

Use Implicit Function theorem to show that \(F(x,y)= y^2+2xy +x-1=0\) can be solved for y in terms of x near all points \((x_0,y_0)\) with \(y_0\neq x_0\). Check the result directly. Also determine if the equation F(x,y) =0 is locally solvable near any point \(x_0,-x_0)\) Please help correcting my solution.

OpenStudy (loser66):

To consider F can be solved for y in terms of x, we need partial \(F_y\neq 0\) . And we got it, \(F_y = 2y-2x \neq 0\)

OpenStudy (freckles):

2y+2x?

OpenStudy (loser66):

To check the result directly, I replace and get \(y=-x\pm \sqrt{x^2-x+1}\) but my prof commented :"This is not the function of x" How to fix?

OpenStudy (loser66):

oh yea, 2y + 2x

OpenStudy (freckles):

because of the +/-

OpenStudy (freckles):

if it was just + then it would be a function or if it was just - then it would be a function but it can't be both like for each x we want one unique y but for each x here we would have at max 2 y

OpenStudy (loser66):

So? I just separate them ?

OpenStudy (loser66):

For the last one, I find \(F(x_0, -x_0)= x_0^2-x_0+1\) and it has no solution. Hence, \(F(x,y) \)is not solvable near any point \((x_0,-x_0)\) . My prof's comment: " Lucky you" It means it is not correct, just lucky. I don't want that luck. I need logic. Please, correct me

OpenStudy (freckles):

can I asked how you got the first relation?

OpenStudy (freckles):

that is the relation $y=-x\pm \sqrt{x^2-x+1}$

OpenStudy (freckles):

\[y=-x\pm \sqrt{x^2-x+1}\]

OpenStudy (loser66):

Just solve F(x,y)=0, that is \(y^2 +2xy +x-1 =0\)

OpenStudy (freckles):

oops just found out

OpenStudy (freckles):

I worked backwards still trying to familiarize myself with the theorem

OpenStudy (loser66):

\(y^2 +2xy+x^2-x^2+x-1=0\\(y+x)^2 =x^2+x-1\)

OpenStudy (freckles):

\[y^2+2xy+x^2-x^2+x-1=0 \\ (x+y)^2=x^2-x+1 \\ \text{ now we have } y \neq -x \\ \text{ so \let's assume } y>-x \text{ then } y+x>0 \\ \text{ so } x+y=\sqrt{x^2-x+1} \\ y=-x +\sqrt{x^2-x+1} \text{ when } y>-x \\ \text{ now } \\ \text{ assume } y<-x \text{ then } y+x<0 \text{ so we have } \\ x+y=-\sqrt{x^2-x+1} \\ y=-x-\sqrt{x^2-x+1} \]

OpenStudy (freckles):

does that make since to you?

OpenStudy (freckles):

sense*

OpenStudy (loser66):

Yes, it is

OpenStudy (freckles):

now the second question you say I kinda don't get it yet are you asking to solve x^2-x+1=0? or is something else meant by solution/

OpenStudy (loser66):

That is when \(y_0= -x_0\), we have \(F(x_0, -x_0)= (-x_0)^2 + 2 (x_0)(-x_0)+x_0-1=0\\-x_0^2+x_0-1=0\)

OpenStudy (freckles):

oh we are dealing with same function

OpenStudy (loser66):

And this equation has no solution, that is at \((x_0,-x_0)\), \(F(x,y) \neq 0\), Hence it is not solvable near there

OpenStudy (freckles):

i mean relation \[F(x,y)=x^2+2xy+x-1\] \[F(x,-x)=x^2+2x(-x)+x-1 \\ F(x,-x)=x^2-2x^2+x-1 \\ F(x,-x)=-x^2+x-1\] so yes I see how you got this

OpenStudy (freckles):

\[x^2-x+1=0 \\ x=\frac{1 \pm \sqrt{1 -4(1)}}{2}\] so you don't want to include imaginary ( I mean complex) solutions?

OpenStudy (loser66):

I think we are working on R^2, not C

OpenStudy (freckles):

I don't know about that one by the way did you asked this on mathexhange someone had your exact same solutions

OpenStudy (loser66):

Yes, :)

OpenStudy (freckles):

I didn't really a straight forward answer for the second question it seemed like they were agreeing with you

OpenStudy (loser66):

I make a lot of question there since from here, not many people help me. :)

OpenStudy (loser66):

My prof took of 5 credits out of 10 for that mistake. I ..... frustrated.

OpenStudy (freckles):

what would happen if we looked at dF/dx instead of dF/dy and solved for x instead

OpenStudy (loser66):

And I got mad with his comment : Lucky you, Ha!!

OpenStudy (loser66):

Nope, I typed it right.

OpenStudy (freckles):

lol made he was trying to bring humor to it all :p

OpenStudy (freckles):

\[F(x,y)=y^2+2xy+x-1=0 \\ \frac{dF}{dx}=2y+1 \neq 0 \\ \text{ means } y \neq \frac{-1}{2} \\ y^2+2xy+x-1=0 \\ y^2+x(2y+1)-1=0 \\ y^2-1=-x(2y+1) \\ x=\frac{y^2-1}{-(2y+1)} \\ \] now what would happen if we use this for our y=-x \[x=\frac{x^2-1}{-(-2x+1)} \\ -x(-2x+1)=x^2-1 \\ 2x^2-x=x^2-1 \\ x^2-x+1=0\] same equation :(

OpenStudy (freckles):

which makes sense I guess

OpenStudy (loser66):

Ok, I will come to see him tomorrow. I need make it clear for my upcoming final. Thanks for the help. I have a bunch of thing need check. Willing to help?

OpenStudy (freckles):

I can only try. I can not guarantee I will be able to help. You have really exceeded my knowledge but I will look at it.

OpenStudy (freckles):

@Michele_Laino we are still trying to figure out why @Loser66 reply to F(x,y)=0 isn't solvable for (x,-x) is wrong

OpenStudy (freckles):

Or I'm still trying to figure it out :p